OpenStudy (anonymous):

simplify logbase6 1/216 please show steps

OpenStudy (ash2326):

\[1/216= 1/6^3\] or \(6^{-3}\) \[ log (6)^{-3}\] \[ -3 log (6)\] =-3

jhonyy9 (jhonyy9):

so because 216=6*6*6 = 6 on exponent 3 - and 1/216 we will get from 6 on exponent minus 3 - so from these resulted that log(base6) 1/216 = log(base6) 6 on exponent minus 3 - and we know that log(base a) a = 1 - so hence we can getting this equal (-3)

jhonyy9 (jhonyy9):

hope so much that is understandably sure